### 03 Feb 2024 v1

# Online Transfer Learning for RSV Case Detection

Abstract
Transfer learning has become a pivotal technique in machine learning, renowned for its effectiveness in various real-world applications. However, a significant challenge arises when applying this approach to sequential epidemiological data, often characterized by a scarcity of labeled information. To address this challenge, we introduce Predictive Volume-Adaptive Weighting (PVAW), a novel online multi-source transfer learning method. PVAW innovatively implements a dynamic weighting mechanism within an ensemble model, allowing for the automatic adjustment of weights based on the relevance and contribution of each source and target model. We demonstrate the effectiveness of PVAW through its application in analyzing Respiratory Syncytial Virus (RSV) data, collected over multiple seasons at the University of Pittsburgh Medical Center. Our method showcases significant improvements in model performance over existing baselines, highlighting the potential of online transfer learning in handling complex, sequential data. This study not only underscores the adaptability and sophistication of transfer learning in healthcare but also sets a new direction for future research in creating advanced predictive models.
Index Terms:
online transfer learning, ensemble method, dynamic weighting mechanism, electronic health record, RSV case detection
I Introduction
Machine learning has made substantial advancements in recent decades, with its applications spanning a wide range of fields such as image and speech recognition, natural language processing, and autonomous driving. Despite these achievements, machine learning in biomedicine faces significant challenges, particularly in data collection. The acquisition of labeled data can be very costly or even unfeasible due to factors like ethical considerations, patient privacy, and the scarcity of certain diseases. These challenges have led researchers to increasingly rely on utilizing data from related domains that have a more abundant supply of data.
To better leverage knowledge from these related yet distinct domains, transfer learning has emerged as a promising approach to enhance model performance. Fundamental to transfer learning are the concepts of the ‘source domain’ and ‘target domain’. The source domain, usually rich in data, serves as the foundation for knowledge extraction. It is characterized by its abundance of well-labeled datasets, which aid models in understanding and encoding complex patterns. In contrast, the target domain, where this pre-learned knowledge is applied, often features a more limited dataset. Transfer learning bridges this gap by adapting models trained on the source domain to function effectively in the target domain, thereby helping address the challenges associated with data scarcity.
Traditional transfer learning approaches [1, 2, 3] typically depend on having complete datasets from both the source and target domains at the start of the training process. However, this is often not feasible, especially in dynamic environments where target domain data is collected progressively. To tackle this issue, the concept of online transfer learning [4, 5, 6] has been introduced. This method is designed for scenarios where target domain data is acquired incrementally over time, while source domain data is already accessible before training begins. In such cases, transferring knowledge from the source domain becomes crucial, particularly because the limited initial data in the target domain may be insufficient for effective learning. The extensive and diverse information available from the source domains can significantly compensate for this shortfall, providing a foundational knowledge base that the model can build upon as more target domain data becomes available. Therefore, the efficiency and effectiveness of learning in the target domain are greatly enhanced by the transferred knowledge from the source domains.
Online transfer learning entails leveraging knowledge from a static source domain and applying it to an ongoing, evolving target task. The model is incrementally updated as new data are received. A crucial element of this process is the utilization of a loss function to fine-tune the model in accordance with the latest incoming data, thus ensuring its continued relevance and efficiency. This adaptability renders online transfer learning particularly suitable for real-time applications, effectively handling issues such as concept drift and providing a more versatile approach to knowledge transfer.
In traditional transfer learning, domains are typically categorized based on similar samples, such as distinguishing between images of cats and dogs, which allows for the straightforward grouping of domains with analogous characteristics. However, online transfer learning presents a nuanced challenge as it focuses on the evolution of a single task over time, making the determination of where to apply temporal segmentation less clear. This difficulty arises because the task itself remains constant, but the data representing the task may change significantly due to temporal factors. This complexity is compounded by the inherent challenge of understanding the relationships among temporally segmented domains. Unlike static categorizations, where similarities and differences are often clear and measurable, the dynamic nature of temporal domains means that the relevance of past data to the present task can be fluid and elusive. For instance, in healthcare, temporal segmentation might involve dividing patient data into annual periods to capture seasonal disease trends. With multiple source domains, integrating them optimally into an ensemble model is challenging. A critical question arises: how do we determine which source domains are most relevant to the current period (or target domain) and which are less relevant or unrelated?
Our proposed method employs a sequential data acquisition strategy to tackle this challenge. Consider a healthcare scenario where the target domain is predicting patient outcomes in a new medical facility. To address initial data scarcity in this facility, our approach significantly relies on pre-trained models derived from source domain data, such as historical patient outcomes from established facilities. These pre-trained source models form the foundation of our ensemble model. This ensemble is adept at autonomously evaluating and determining the relevance of each source model’s learned patterns to the current patient data in the new facility. For instance, a source model trained on a large urban hospital might be initially prioritized for its broad spectrum of cases. As more patient data accumulates in the new facility, the ensemble dynamically adjusts its reliance on this urban model, potentially shifting focus to another source model that aligns more closely with emerging local trends. This strategic utilization of pre-trained source models ensures that the most effective ones are prioritized during the critical early stages of training in the target domain.
As the target domain accumulates more data, the ensemble model modifies its approach. Initially dependent on pre-trained source models, it gradually begins to prioritize the continually updated target model. This change is twofold: a shift in reliance from source to target models and a nuanced adjustment in the weight each model holds in decision-making. This transition enables the model to evolve alongside the growing target domain dataset. The distinct roles of shifting focus and adjusting weights demonstrate the adaptability of our online transfer learning approach, particularly in environments where data is collected over time.
Our framework makes two main contributions. First, we introduce an online multi-source transfer learning architecture to study biomedical information over multiple seasons. Second, we explore the performance of different ways of generating weights for the ensemble model to improve its accuracy.
We evaluate the effectiveness of our approach using real-world Respiratory Syncytial Virus (RSV) data that were collected over multiple seasons from patient encounters in Emergency Departments (EDs) at the University of Pittsburgh Medical Center (UPMC). Our primary goal is to utilize reported signs and symptoms to accurately classify patients as either having RSV or not, a critical step towards improving diagnosis and treatment strategies. RSV is a major cause of respiratory infections that lead to hospitalizations [7]. The burden of RSV continues to grow, especially in young children aged less than 5 years old[8, 9]. It has been also identified as an important disease causing hospitalizations of the elderly and high-risk adults [10].
Characteristically, RSV exhibits a seasonal pattern in case numbers [11]. In our approach, we specifically segment the RSV case data into distinct domains based on a defined annual cycle, with each domain spanning from June of the previous year to May of the current year. This segmentation strategy is crucial to our framework, aiming to enhance the predictive accuracy of RSV case detection by incorporating data from multiple preceding seasons as source inputs. For instance, in our experiments, we use patient encounter data ranging from June 2011 to May 2018 to develop an ensemble model capable of classifying encounters in the period from June 2018 to May 2019. The experimental results support that our approach improves predictive performance, particularly related to public health planning and resource allocation for anticipated RSV seasons.
II Related Work
Traditional transfer learning methodologies typically operate under the premise that the source and target domains are predefined, with the primary objective being the transference of knowledge from the source to the target domain. Numerous comprehensive surveys [12, 13, 14, 15, 16, 17] have thoroughly examined and classified a broad range of transfer learning approaches, considering factors like algorithm types and their various use cases. A notable recent development is the growing body of literature reviews focusing on the integration of transfer learning with other advanced machine learning techniques. These include combinations with deep learning [18, 19], reinforcement learning [20, 21, 22], and federated learning [23, 24]. Additionally, some reviews have specifically targeted the application of transfer learning in key areas such as computer vision (CV) [25, 26, 27, 28], natural language processing (NLP) [29, 30, 31], medical image analysis [32, 33], and wireless communication [34, 35], providing insights into the practical implications and utility of transfer learning in these specific fields.
Online transfer learning has emerged to address situations where data are received sequentially and are not available at the start of training. This approach is particularly relevant in dynamic environments where real-time data processing is crucial. Some popular online transfer learning methods have demonstrated advanced performance in many real-world scenarios. Online transfer learning methods can be categorized into data-based methods and model-based methods.
Data-based transfer learning methods involve the transfer of either the original data or their transformed features to a target domain, allowing the target models to be trained with these augmented data, thereby enriching the available data within the target domain. Online Heterogeneous Knowledge Transition (OHKT) [36] initialize the target model using co-occurrence data with pseudo labels generated from the source model.
Conversely, model-based transfer learning methods focus on building models within the source domain, adapting both the model structure and parameters from this domain. Here, ‘model structure’ refers to the architecture or configuration of the model (such as layers in a neural network), while ‘model parameters’ are the specific weights and biases that are learned from the source data. The target models utilize these insights from the source domain, with the aim of potentially enhancing performance. Online Heterogeneous Transfer by Hedge Ensemble (OHTHE) [37] handles the online heterogeneous transfer problem by updating weights of offline and online classifiers based on the Hedge weighting strategy.
For our online transfer learning algorithm, we utilized RSV detection as an illustrative example task. Machine learning algorithms have been increasingly utilized for detecting RSV. Tso et al. [38] developed an extreme gradient boost decision tree model to detect RSV using clinical measurements in electronic health records (EHR), such as respiratory rate and peripheral oxygen saturation, from EDs and inpatient visits in 2019. In a separate study, Cord et al. [39] evaluated various methods, including Logistic Regression, Gaussian Naive Bayes, and Random Forest, before selecting the Light Gradient Boosting Machine algorithm for their RSV detection models.
III Methodology
In this section, we begin by outlining the problem formulation for the online multiple source-domain scenario. Following this, we delve into the base model, setting the stage for the introduction of our proposed method.
III-A Problem Formulation
In this study, we address the challenge of binary classification in environments where data are not static but arrive sequentially, often correlated with specific time frames such as seasons. Our focus is on developing a method for studying information collected over seasons using Bayesian networks as the foundation of our classifier training.
Bayesian networks, a type of probabilistic graphical model, are particularly well-suited for this task. They express conditional dependencies between a set of variables through a Directed Acyclic Graph (DAG). In these networks, each node represents a variable, encompassing both input features $\mathbf{x}\in\mathbb{R}^{d}$ and binary labels $y\in\{0,1\}$. The edges denote probabilistic dependencies between these variables. This structure allows Bayesian networks to model complex relationships within the data and to infer the likelihood of certain outcomes given a set of observed variables, thereby making them ideal for classification tasks.
Given a sufficient number of training instances $(\mathbf{x},y)$, Bayesian networks are adept at predicting the probability that a known cause (represented by the feature vector $\mathbf{x}$ was the contributing factor for the observed event (the label $y$). This probabilistic approach not only offers predictions but also provides insights into the strength and nature of the relationships between different features and the outcome.
In real-world applications, each epidemic season’s data can be considered a distinct domain. This temporal characteristic poses a unique challenge, particularly at the start of a new period when data availability is limited. To address this, we propose a transfer learning approach where the current season is treated as the target domain, and data from multiple previous seasons are utilized as source domains. Our method employs data from these multiple source domains, denoted as $D^{S}=\{D_{S}^{1},D_{S}^{2},\ldots,D_{S}^{n}\}$, where each $D_{S}^{i}$ corresponds to a specific previous season’s data, alongside a single target domain $D_{T}$. We assume that patterns and knowledge from previous seasons can be leveraged to effectively predict or analyze trends in the current season, despite the initial scarcity of data. The primary goal is to leverage historical data to inform and improve the predictive model for the current season, under the assumption that underlying patterns and relationships in the data persist over time.
III-B Base Model
Our methodology initiates with the implementation of the Naive Bayes classifier, a cornerstone of probabilistic classification techniques. Central to Naive Bayes is the simplification of probability calculations, achieved by the assumption of conditional independence among features given a class label. This assumption allows the classifier to compute the probability of a class label $y$ given a set of features $x_{1},\ldots,x_{n}$ using the formula:
$P\left(y\mid x_{1},\ldots,x_{n}\right)$ $=\frac{P(y)P\left(x_{1},\ldots,x_{n}\mid y\right)}{P\left(x_{1},\ldots,x_{n}\right)}$ (1)
$=\frac{P(y)\prod_{i=1}^{n}P\left(x_{i}\mid y\right)}{P\left(x_{1},\ldots,x_{n}\right)}.$ (2)
The effectiveness of Naive Bayes lies in its simplicity and its computational and statistical efficiency, especially in cases where the features are reasonably independent given the class label.
III-C Online Multi-source Ensemble Model
In our framework, for each source domain $D_{S}^{i}$, we train a unique offline base model $f_{S}^{i}$ as a classifier on this domain. Similarly, a specific online training classifier $f_{T}$ is designated for the target domain. The architecture of our proposed model is depicted in Fig. 1. Central to our approach is an ensemble model, formulated as:
$f=w_{1}f_{S}^{1}+\ldots+w_{n}f_{S}^{n}+w_{T}f_{T}$ (3)
Here, $w_{i}$ represents the weight parameter for the $i$-th source domain classifier, and $w_{T}$ denotes the weight parameter for the target domain classifier. Crucially, each source domain classifier $f_{S}^{i}$ is pre-trained on its respective domain data, while the target domain classifier $f_{T}$ is dynamically trained and updated with new instances from the target domain as they emerge.
Figure 1: Architecture of the proposed ensemble model.
Our ensemble method capitalizes on the strengths of individual classifiers trained on distinct historical data patterns from each source domain, integrating these with the versatile classifier of the target domain. This integration aims to enhance the model’s overall predictive capabilities and accuracy, which is especially beneficial when dealing with limited initial data in the target domain. For weight allocation in our ensemble model, adaptive weights provide a dynamic adjustment mechanism, encompassing two strategies:
Predictive Performance-Adaptive Weighting
This Predictive Performance-Adaptive Weighting (PPAW) strategy involves updating all weights, including those of the target domain classifier and source domain classifiers, solely based on prediction outcomes as shown in Algorithm 1. As new training instances are processed, the model evaluates the accuracy of each classifier, adjusting weights based on their predictive performance: weights of accurate classifiers remain the same while inaccurate classifiers receive a weight penalty, reducing their influence in future predictions.
Input: Pre-trained source models $f_{S}^{1}(\cdot),\ldots,f_{S}^{n}(\cdot)$, Target model $f_{T}(\cdot)$, Target Dataset $X=\{X_{j}\}$ for $j=1,2,\ldots$, Weight penalty $\alpha\in\left(0,1\right)$, Target weight factor $\beta\in\left(0,1\right)$. Initialization: Initialize source model weights $w_{S}^{i}=1.0$ for $i=1,2,\ldots,n$ and target model weight $w_{T}=1.0$. while receive new data point $X_{j}$ do
       if algorithm = PVAW then
             Adjust target weight: $w_{T}\leftarrow w_{T}\times\beta\times j$
       end if
      Normalize weights:
            $w_{S}^{i}\leftarrow\frac{w_{S}^{i}}{\sum_{i=1}^{n}w_{S}^{i}+w_{T}},$       $w_{T}\leftarrow\frac{w_{T}}{\sum_{i=1}^{n}w_{S}^{i}+w_{T}}$
       Calculate combined probability for final prediction:
            $prob=\sum_{i=1}^{n}w_{S}^{i}\times f_{S}^{i}(X_{j})+w_{T}\times f_{T}(X_{j})$ for $i=1$ to $n$ do
             if $f_{S}^{i}(X_{j})$ predicts incorrectly then
                   Adjust weight: $w_{S}^{i}\leftarrow\alpha\times w_{S}^{i}$
             end if
       end for
      if $f_{T}(X_{j})$ predicts incorrectly then
             Adjust weight: $w_{T}\leftarrow\alpha\times w_{T}$
       end if
      Update $f_{T}$ with $X_{j}$;
       Return $prob$ as the predicted probability for $X_{j}$;
end while
Algorithm 1 Predictive Adaptive Weighting Strategies: Performance-Adaptive (PPAW) and Volume-Adaptive (PVAW)
Predictive Volume-Adaptive Weighting
The Predictive Volume-Adaptive Weighting (PVAW) strategy advances beyond the PPAW by innovatively adjusting the weights of classifiers in response to changes in data volume and prediction accuracy. For the target domain classifier, its weight is dynamically modified based on two key factors: the growing volume of target data and its prediction results. As the amount of target domain data increases, the classifier’s weight is proportionally enhanced. This increase is quantified by multiplying the weight with a factor that signifies the quantity of target data, a process facilitated by the variable $j$. The variable $j$ serves as a sequential index for each new data point, effectively tracking the expansion of the target dataset. As more data points are received (i.e., as $j$ increases), the target classifier’s weight is correspondingly adjusted upwards. This adjustment ensures that the weight of the target classifier is attuned to the scale of the available data, enabling the model to progressively prioritize more recent and relevant insights. Simultaneously, the source domain classifiers’ weights are adjusted primarily based on their prediction performance. This dual mechanism—adapting weights for both prediction accuracy and data volume—enables a more dynamic and responsive approach in the PVAW strategy. The target classifier’s weight becomes increasingly significant as more target data becomes available, ensuring that the model’s predictions are continually refined and aligned with the latest trends and patterns in the target domain.
The PPAW and PVAW strategies, detailed in Algorithm 1, aim to optimize the ensemble model’s performance. PPAW modifies response based on classifier accuracy, while PVAW further tailors the target model’s weights to the growing data volume in the target domain. This method ensures the ensemble model stays attuned to the target domain’s evolving dynamics, prioritizing classifiers that resonate with the latest data trends.
We hypothesize that the adaptive nature of learnable weights will demonstrate superior performance compared to static weight assignment strategies. This hypothesis is predicated on the notion that dynamically adjusting weights will allow the ensemble model to evolve in sync with the incoming target data, potentially leading to more accurate and adaptable predictions.
IV Dataset
The research data were sourced from the University of Pittsburgh’s Neptune Research Data Warehouse [40], and the study protocol received approval from the University of Pittsburgh Institutional Review Board (Protocol STUDY20030193). The dataset consists of 2,263,360 encounters from five EDs at the UPMC over nine years from January 2011 to May 2020. In regions with temperate climates, RSV infections predominantly manifest during the colder winter months. The onset of these infections is usually observed around October and November in the Northern Hemisphere. Typically, the highest incidence of RSV cases, or the peak period, occurs in January and February. As the season progresses, the frequency of RSV infections starts to decline, with the final cases of the season generally being recorded between March and May [11]. Given the seasonal nature of RSV outbreaks, we defined one season as spanning from June of one year to May of the following year, rather than using a calendar year division for data segmentation. Table I details the sizes of encounters for each season. An encounter was classified as related to RSV if it included one of any of the following International Classification of Diseases (ICD) codes at discharge:
  • 79.6 Respiratory syncytial virus
  • 466.11 Acute bronchiolitis due to respiratory syncytial virus
  • 480.1 Pneumonia due to respiratory syncytial virus
  • B97.4 Respiratory syncytial virus as the cause of diseases classified elsewhere
  • J12.1 Respiratory syncytial virus pneumonia
  • J20.5 Acute bronchitis due to respiratory syncytial virus
  • J21.0 Acute bronchiolitis due to respiratory syncytial virus
Encounters that did not include any of these ICD codes were categorized as non-RSV.
Our dataset contains 7614 features, including demographic information, lab test results, and natural language processing extracted clinical findings, detailed in Table II. The clinical findings were obtained by using the MetaMap parser [41], which applies pattern-matching and deduction rules. The clinical notes for each encounter serve as input for the parser, which extracts a set of clinical findings and their values. The values for each finding are either present (P), negated (N), or missing (M), where missing indicates that clinicians did not mention the finding in the note, according to the parser. In cases where an encounter had multiple clinical notes, we used the union of all clinical findings across all the notes. Contradictions across notes were resolved in favor of present findings, and this also applied to multiple occurrences of a finding in a single note.
TABLE I: Sizes of datasets
Season Number of RSV Labeled Cases Number of Non-RSV Labeled Cases
2011-2012 251 259,596
2012-2013 238 258,024
2013-2014 132 254,076
2014-2015 179 250,476
2015-2016 200 252,432
2016-2017 381 257,172
2017-2018 408 243,708
2018-2019 507 240,312
2019-2020 888 244,380
TABLE II: Features in research datasets
Features Distinct Values (Descriptions)
age
less5 ($<$5), ge5less18 (5-17), ge18less65 (8-64), ge65 (65+), M (value is missing)
gender
female, male, unknown, M (value is missing)
ethnicity
Hispanic or Latino, not Hispanic or Latino, not specified, M (value is missing)
race
black, white, other, not specified, M (value is missing)
7541 UMLS CUIs extracted from free-text reports by MetaMap
P (present), N (negated), and M (value is missing). One CUI example is C0859927 (Increased work of breathing).
69 laboratory results
H (high), L (low), N (normal), M (value is missing). One lab example is oxygen saturation.
Figure 2: Change in weights over time.
Note that the x-axis is scaled non-uniformly to enhance data visibility.
Figure 2: Change in weights over time.
V Experiments & Results
V-A Experiment Settings
In this study, we employed an online multi-source transfer learning ensemble model to enhance our model’s adaptability in a specific target domain. This model uniquely incorporated datasets from previous seasons, each treated as a distinct source domain. Specifically, for targeting a particular season, such as 2019-2020, the model was trained on data from individual seasons preceding 2019-2020, with each season (2011 to 2018) considered separately. This approach ensures that each season’s unique data characteristics are captured effectively.
This approach mirrors real-world situations, where models must adjust to new and varied data from the target domain while leveraging knowledge from historical data. This is particularly useful in biomedical fields where labeled data may be limited, thereby improving the model’s ability to generalize and maintain robustness.
To more closely replicate real-world conditions, our model was set in an online learning environment for the target domain, updating continually with each new data point. This mimics the ongoing influx of medical data in actual scenarios, ensuring our model remains current and effective.
We evaluated our model’s performance using the Area Under the Receiver Operating Characteristic Curve (AUROC), an important measure for gauging a model’s discriminative performance when classifying positive and negative cases across diverse data sequences. This offers a comprehensive assessment of the model’s overall effectiveness.
In the specified algorithm, two hyperparameters, $\alpha$ and $\beta$, play crucial roles in the dynamic adjustment of model weights during the learning process. The setting of $\alpha$ and $\beta$ is consistent across all domains. The hyperparameter $\alpha$ is set to $\log_{2}10$ = 3.32, a value carefully selected after experimenting with several hyperparameter settings. This medium value for $\alpha$ was chosen to balance the weight penalty applied to a classifier’s weight upon making an incorrect prediction. This particular value for $\alpha$ strikes a balance between being too lenient and too punitive, allowing for a gradual decrease in a classifier’s influence while still retaining its contributions where they prove useful.
The hyperparameter $\beta$, set at $1/200000$, plays a critical role in the volume-adaptive component of our algorithm. This parameter is instrumental in fine-tuning the adjustment of the target weight $w_{T}$, aligning it with the growing volume of incoming data. Here, the factor $j$ serves as an index representing the sequence of data points being processed. The specific choice of $1/200000$ for $\beta$ strikes a delicate balance. When the number of instances has not yet reached 200,000, this configuration ensures that the target weight is subtly suppressed, preventing premature dominance in the model’s decision-making process. However, once the count surpasses this threshold, the target weight is progressively amplified, reflecting the increased reliability and relevance of the more substantial data corpus. This mechanism ensures that the target model’s influence is appropriately scaled, growing more pronounced as the volume of target data becomes more significant, thereby sensitively responding to the increase in data volume while simultaneously tempering the rate of weight adjustment.
By adopting this value for $\beta$, the algorithm ensures a balanced and responsive adaptation to the accumulating target data. Such a setup allows our ensemble model to systematically transition its focus, increasingly prioritizing the most recent data, which is presumed to be the most relevant for the current learning context. This approach exemplifies the nuanced and progressive nature of online learning, where the model dynamically adapts to the evolving data landscape, ensuring that its predictions remain accurate and pertinent.
V-B Results and Analysis
Our primary objective is to develop a model adept at classifying data from the current season, which is data from the season 2019-2020 in this study. To achieve this, we trained individual models for each season ranging from 2011 to 2018. Subsequently, we evaluated the performance of these models using data from the season 2019-2020 to ascertain their effectiveness in handling present-season data. The results of these assessments are comprehensively tabulated and presented in Table III, providing a clear overview of each model’s performance across this temporal span.
TABLE III: Results of individual models
Season AUROC
2011-2012 0.800
2012-2013 0.628
2013-2014 0.637
2014-2015 0.765
2015-2016 0.838
2016-2017 0.867
2017-2018 0.866
2018-2019 0.865
For the results, a clear upward trend in AUROC scores is evident as we approach the season of 2019-2020. This trend is especially pronounced in models trained with more recent data, indicating an improvement in the classifiers’ ability to distinguish between positive and negative cases as they are trained on data closer to the target season. This improvement can be attributed to the increasing relevance and recency of the data. As the training data gets closer to the target season, it likely becomes more representative of the current patterns and trends, thereby enhancing the classifiers’ predictive accuracy.
To evaluate the performance of PPAW and PVAW strategies, we compare them with baseline methods, including:
  • Online Learning Model: This model employs an online learning strategy, focusing exclusively on sequentially acquired data from the target domain for training.
  • Model Trained from All Previous Data: This approach involves pre-training a model using a combined dataset from all source domains.
  • Ensemble Model with Equal Weights: This method employs an ensemble model that integrates various individual source models, assigning them equal and fixed weights throughout the process.
  • Ensemble Model with Volume-based Weights: This strategy uses an ensemble model that weights individual source models according to the volume of their respective training datasets.
  • Ensemble Model with Time-based Weights: This ensemble method integrates individual models based on their temporal proximity to the testing season (target domain), defining ‘time distance’ as the number of seasons separating a model’s training data from the testing period.
The final three approaches within the ensemble model framework employ predetermined weights, which are determined in one of three ways: evenly distributed across all models, adjusted according to the sizes of their training datasets, or allocated baspleed on the models’ temporal proximity to the target domain’s data period.
In our comparative analysis of various online learning strategies against the backdrop of sequential epidemiological data, distinct performance metrics emerge, highlighting the efficacy and adaptability of these models as shown in Table IV. The Online Learning Model, with an AUROC of 0.850 (P$=$0.037), showcases a commendable capacity for leveraging sequentially acquired data, underscoring the viability of online learning in dynamically evolving domains. Slightly surpassing this, the Model Trained from All Previous Data, at an AUROC of 0.853 (P$<$0.001), demonstrates the potential advantages of a comprehensive, cross-domain dataset for pre-training, affirming the value of extensive historical data. Conversely, the Ensemble with Equal Weights method, despite its statistical significance, registers a lower AUROC of 0.728 (P$<$0.001), suggesting that uniform weighting across diverse models may dilute the predictive precision. More notably, the Ensemble with Volume-based Weights, at an AUROC of 0.680 (P$<$0.001), indicates that volume-centric model weighting may not effectively capture the intricate data nuances necessary for accurate target domain predictions. However, an improvement is observed with the Ensemble with Time-based Weights, achieving an AUROC of 0.761 (P$<$0.001), which implies that temporal relevance plays a crucial role in enhancing model performance, albeit not to the extent of the more sophisticated methods. Remarkably, both the Predictive Performance-Adaptive Weighting (PPAW) and the Predictive Volume-Adaptive Weighting (PVAW) strategies exhibit superior performance metrics, with PPAW matching the Model Trained from All Previous Data at an AUROC of 0.853 (P$<$0.001), and PVAW leading the cohort with an AUROC of 0.870. This delineates the PVAW strategy’s predominant advantage in dynamically calibrating model weights based on the evolving data landscape, marking a significant stride towards optimizing predictive models for healthcare applications, particularly in the nuanced domain of Respiratory Syncytial Virus (RSV) data analysis.
TABLE IV: Comparative results of different models, featuring P values from DeLong’s test for statistical differences between their respective AUROCs and the AUROC of the PVAW method.
Method AUROC
Online Learning 0.850 (P=0.037)
Trained from All Previous Data 0.852 (P$<$0.001)
Ensemble with Equal Weights 0.728 (P$<$0.001)
Ensemble with Volume-based Weights 0.680 (P$<$0.001)
Ensemble with Time-based Weights 0.761 (P$<$0.001)
PPAW 0.853 (P$<$0.001)
PVAW 0.870
Figure 3: Possibility of features present given RSV.
Figure 4: $Log_{10}LR$ of features present.
Our proposed PVAW approach, which utilizes dynamically updated weights, surpasses these conventional baseline methods. The essence of its enhanced performance lies in its ability to dynamically adapt, customizing its relevance to the current data. This dynamic adaptation is achieved by continuously adjusting the weights of the models based on their immediate performance with new data. Such a strategy enables the model to respond effectively to changing data trends and to identify patterns that were not apparent in earlier training datasets. As a result, our method shows a marked improvement in both accuracy and effectiveness in classifying medical data, outstripping the other methodologies evaluated in the study.
Fig. 2 illustrates the dynamic changes in the model weights corresponding to data from different seasons as new instances are introduced over time, providing a visual correlation with the tabulated performance metrics. Each line on the graph represents the weight trajectory for a model trained on data from a specific season, spanning from 2011 to 2018. Initially, all model weights start with a high value, denoting an equal influence within the ensemble. As the number of instances increases, the graph shows significant fluctuations in these weights. Notably, models corresponding to 2016-2017 and 2017-2018 maintain higher weights throughout, signifying their sustained predictive relevance over time. This trend aligns with Table III’s results, where the AUROC scores for 2016-2017 and 2017-2018 are the highest among the evaluated seasons. These scores indicate that the classifiers trained on 2016-2017 and 2017-2018 data are most effective at accurately predicting outcomes for the target season of 2019-2020. The congruence between the graph’s weight trends and the AUROC scores in Table III emphasizes the ensemble model’s capability to adaptively prioritize the most relevant historical data for current predictions, enhancing the overall accuracy of the RSV case detection model.
Fig. 3 and Fig. 4 present the possibility of features present in the RSV positive cases and the $Log_{10}$ likelihood ratio (LR) of variates respectively. 13 variates with the highest LR (with possibility$>$0.01) in our target season 2019-2020 were selected as the comparative features. In Fig. 3, the possibility of Age - less than 5 appears high with values ranging from 0.68 to 0.96 in all years, while other features only have around or below 0.2 for possibility. As shown in Fig. 4, with relatively small ranges of values, Age - less than 5 and Increased work of breathing are two much more steady features than others observed through all years. Compared to season 2019, the data points in seasons 2012-2013 and 2013-2014 have been the farthest from the 2019-2020 season, suggesting that the feature patterns in the 2012-2013 and 2013-2014 seasons are the least similar to our target season in 2019-2020, which also aligns with our model results in Table III that seasons 2012-2013 and 2013-2014 present the lowest AUROC. In contrast, features from seasons 2017-2018 and 2018-2019 are mainly located closest to season 2019-2020, and their individual models also provide good performance as shown in Table III.
VI Discussion
In this study, we have analyzed various ensemble methods applied to the challenging task of medical data classification. Our primary objective was to develop a robust classifier for the season of 2019-2020, leveraging historical data spanning from 2011 to 2018. The results obtained from our experiments reveal a temporal trend in the performance of individual models. Contrary to the straightforward expectation that more recent data would inherently yield better results, the final weights of the ensemble model exhibited a complex relationship with the season of training data. Notably, the maximum weight was achieved with data from season 2017-2018, highlighting the importance of allowing the model to autonomously discern relevant patterns and relationships. This finding emphasizes the importance of adopting dynamic and data-driven approaches in medical data classification, as the temporal relevance of information may not always align with conventional intuitions.
The evaluation of ensemble methods against baseline models provided valuable insights. Our proposed PVAW method outperformed other techniques. This superior performance can be attributed to the method’s inherent adaptability, which allows it to continuously learn and assign weights to individual models based on their performance on the most recent data. By doing so, our approach effectively captures emerging patterns and evolving relationships within the data, resulting in enhanced classification accuracy. In an environment where medical knowledge and practice are continually evolving, our research underscores the potential of dynamic ensemble methods to provide more accurate and efficient solutions for medical data classification tasks. These findings have significant implications for healthcare decision-making, as they demonstrate the importance of data-driven adaptability in capturing the intricate nuances of medical data.
In spite of the promising results presented in this study, there remain several noteworthy limitations that warrant attention in future research. In our study, while the Naive Bayes models proved effective, they represent only a portion of the potential modeling techniques available for medical data classification. These models, although robust, may not fully capture the complexities of medical data as efficiently as more advanced algorithms, such as neural networks. Therefore, future research should explore incorporating these advanced machine learning techniques, which may offer an improvement in the modeling of the nuanced patterns in medical data. Additionally, the dynamic penalty factor used in our study, though effective, has room for refinement. For example, enhancing its adaptability to the changing data distributions in the target domain could lead to more accurate and responsive models.
Another limitation of our study is the specificity of our dataset, which was centered around RSV information from UPMC. This presents a potential constraint in terms of the generalizability of our findings across different medical contexts. The healthcare field is diverse, and patterns observed in one dataset may not be universally applicable. Our future research will aim to validate the model across a broader spectrum of biomedical scenarios, for example, we can incorporate data from different medical institutions. This would help assess the transferability of our approach and determine necessary modifications for different medical settings. Expanding the scope of our research through collaborative and cross-disciplinary studies could also provide new insights and data sources, enhancing the model’s applicability and robustness.
VII Conclusion
In our study, we introduced an innovative online multi-source transfer learning methodology with a dynamic weighting mechanism and applied it to analyzing RSV data, systematically gathered over several seasons from actual patient encounters at UPMC. Our experimental results robustly demonstrate that this method significantly outperforms established baseline techniques, highlighting the considerable potential of our framework. Notably, it adeptly addresses the challenges inherent in biomedical research, such as the scarcity of labeled data and the complexities associated with sequential data. The success of our approach not only affirms its effectiveness but also paves the way for advanced predictive models in healthcare, offering a promising avenue for future explorations in the domain of sequential epidemiological data analysis.
VIII Acknowledgements
This work was supported by the research grants R00LM013383 and R01LM013509 from the National Library of Medicine, National Institutes of Health. The research data were retrieved by the University of Pittsburgh’s Health Record Research Request (R3), which was supported by the National Institutes of Health through the grant UL1 TR001857. Dr. Harry Hochheiser was supported by MIDAS grant U24GM132013. Dr. Jessi Espino was partially supported by the National Center for Immunization and Respiratory Diseases grant 5U01IP001184. We are truly grateful for the invaluable assistance and guidance from Dr. Richard K. Zimmerman and Dr. Mary Patricia Nowalk, epidemiologists at the University of Pittsburgh; Dr. Milos Hauskrecht, a professor of computer science from the University of Pittsburgh.